You follow the lab procedure and at the end of Part 4C, you used an average volume of 25.0 mL of the 0.0180 mol/L titrant to reach the end point; 50.0 mL samples of lake water were used each trial.

Please calculate the Total Alkalinity (eq/L) of this lake water sample.
Answer #1

We know, M1V1 = M2V2

Here 25ml of 0.0180 mol/L titrant is used to reach end point of 50ml sample.

Here, M1= 0.0181 , V1= 25 & M2= ?, V2= 50

Hence concentration of sample is : (M2)

25×0.0180 = M2 × 50

=> M2 = (25×0.0180)/50

=> M2 = 0.009mol/L

​​​